# ESDD2: Environment-Aware Speech and Sound Deepfake Detection Challenge Evaluation Plan

### v2

@@ Abstract
Audio recorded in real‑world environments often contains a mixture of foreground speech and background environmental sounds. With rapid advances in text‑to‑speech, voice conversion, and other generation models, either component can now be modified independently. Such component-level manipulations are harder to detect, as the remaining unaltered component can mislead the systems designed for whole deepfake audio, and they often sound more natural to human listeners. To address this gap, we have proposed CompSpoofV2 dataset [^1] CompSpoofV2: https://xuepingzhang.github.io/CompSpoof-V2-Dataset/ and a separation-enhanced joint learning framework [^2] Baseline: https://github.com/XuepingZhang/ESDD2-Baseline. CompSpoofV2 is a large-scale curated dataset designed for component-level audio anti-spoofing, which contains over 250k audio samples, with a total duration of approximately 283 hours. Based on the CompSpoofV2 and the separation-enhanced joint learning framework, we launch the Environment-Aware Speech and Sound Deepfake Detection Challenge (ESDD2) [^3] ESDD2: https://sites.google.com/view/esdd-challenge/esdd-challenges/esdd-2, focusing on component-level spoofing, where both speech and environmental sounds may be manipulated or synthesized, creating a more challenging and realistic detection scenario. The challenge will be held in conjunction with the IEEE International Conference on Multimedia and Expo 2026 (ICME 2026) [^4] ICME 2026: https://2026.ieeeicme.org/.
@@ I Introduction
In real-world recording scenarios, audio often contains a mixture of foreground speech and background sounds. We define such audio as comprising two components: (i) speech, i.e., linguistically meaningful speech produced by the primary speaker, and (ii) environmental sound, i.e., any non‑speech background or non‑target speech. With rapid advances in text‑to‑speech, voice conversion, and other generation models, either component can now be modified independently. For example, replacing the background while leaving the foreground human speech unchanged, or modifying the speech content while preserving the background. Such component-level manipulations are harder to detect, as the remaining unaltered component can mislead the systems designed for whole deepfake audio, and they often sound more natural to human listeners.
To address this gap, we introduced a component-level spoofing audio dataset CompSpoof [16] in 2025. Then, we extended the dataset to CompSpoofV2, which contains more than 250,000 audio clips (around 283 hours) formed by mixing bona fide and spoofed speech and environmental sound from multiple sources. We also proposed a separation-enhanced joint learning framework to explore to address this gap.
@@ II Dataset
The CompSpoofV2 dataset contains more than 250,000 audio clips (around 283 hours) formed by mixing bona fide and spoofed speech and environmental sound from multiple sources. Each audio belongs to one of five categories, covering all combinations of bona fide and spoofed speech and environmental sound as shown in Table I. The task of this challenge is to classify a given audio clip into one of the five defined classes, as shown in Figure 1.
@@ 4 Spoofed Spoofed spoof_spoof Spoofed speech mixed with spoofed environmental sound
The audio sources for each category are summarized in Table II and Table III.
@@ 4 ASV5, MLAAD, New Generated EnvSDD, VcapAV, New Generated
III Baseline
We propose a separation-enhanced joint learning framework as the baseline for this task. The framework first detects potentially spoofed mixtures, then separates the audio into speech and environmental components, which are then processed by speech- and environment-specific anti-spoofing models. Their outputs are fused and mapped to five-class predictions. The separation and anti-spoofing models are jointly trained to preserve spoofing-relevant cues. More detail can be found in [16]. The baseline results are as IV.
Original EER Speech EER Env. EER F1-score
Val. set 0.0031 0.0172 0.3766 0.9462
Eval set 0.0174 0.1993 0.4336 0.6224
Test set 0.0173 0.1978 0.4279 0.6327
IV Evaluation Criteria
@@
In addition, we introduce three auxiliary metrics based on equal error rate (EER): $\mathrm{EER}_{\text{original}}$, $\mathrm{EER}_{\text{speech}}$ and $\mathrm{EER}_{\text{env}}$. $\mathrm{EER}_{\text{original}}$ is the EER for distinguishes the Original ($i=0$) class from all other four classes ($i=\{1,2,3,4\}$. Similarly, $\mathrm{EER}_{\text{speech}}$ is the EER for detecting spoofed speech component, while $\mathrm{EER}_{\text{env}}$ is the EER for detecting spoofed environmental component. The three EER metrics are included for diagnostic and analysis purposes only; they will not be considered for leaderboard ranking.
@@ V Submission Guideline
We use CodaBench platform for results submission [^5] Results Submission: https://www.codabench.org/competitions/12365/.
Participants are expected to compute and submit a text file. We suppose two submission file formats. One is the two-line format as follows. In the file, each line includes the audio ID and the predicted class ID, which is an integer label from 0 to 4 corresponding to the class ID in Table I.
1 UGBFJ3.wav | 0
2 ABCD12.mp3 | 3
The other is the five-line format as follows. In the file, audio ID, the confidence score for predicting the original class, the confidence score for the speech component, the confidence score for the environmental sound component, and the predicted class ID.
1 UGBFJ3.wav | 1.22 | 0.91 | 0.99 | 0
2 ABCD12.mp3 | -1.54 | 0.89 | -1.89 | 3
The challenge has two phases: the preparation phase and the final phase. During the preparation phase, participants submit evaluation set predictions and the leaderboard reflects evaluation set rankings; and the final evaluation phase, where participants submit test set predictions and the leaderboard determines the final ranking. Each team may submit up to ten results in the final evaluation phase, with the best three used for final ranking to ensure fairness. After the competition, we will release the metadata of the evaluation and test sets to support reproducibility and future research.
@@ VI Rules
  • Publicly available pre-trained models are allowed but must be disclosed.
@@
  • The use of any synthetic audio—including audio generated by text-to-Speech models, voice conversion models, or any other audio generation models—is strictly prohibited.
  • Participants may use additional public datasets only with the organizers’ prior approval. Requests must be submitted before February 20, 2026, and the organizers will publish the list of approved datasets; any datasets not on this list are not permitted for training.
@@ VIII Important Dates
Training, validation and evaluation sets release; Baseline code release; Preparation phase leaderboard update January 30th, 2026
@@
Test set release; Final evaluation phase leaderboard update March 20th, 2026
@@
More information can be found in the Google Group [^7] Google Group: https://groups.google.com/g/esdd2-challenge

### v3

@@ Abstract
Audio recorded in real‑world environments often contains a mixture of components: foreground speech and background environmental sounds. With rapid advances in text‑to‑speech, voice conversion, and other generation models, either component can now be modified independently. Such component-level manipulations are harder to detect, as the remaining unaltered component can mislead the systems designed for whole deepfake audio, and they often sound more natural to human listeners. To address this gap, we have proposed CompSpoofV2 dataset [^1] CompSpoofV2: https://xuepingzhang.github.io/CompSpoof-V2-Dataset/ and a separation-enhanced joint learning framework [^2] Baseline: https://github.com/XuepingZhang/ESDD2-Baseline. CompSpoofV2 is a large-scale curated dataset designed for component-level audio anti-spoofing, which contains over 250k audio samples, with a total duration of approximately 283 hours. Based on the CompSpoofV2 and the separation-enhanced joint learning framework, we launch the Environment-Aware Speech and Sound Deepfake Detection Challenge (ESDD2) [^3] ESDD2: https://sites.google.com/view/esdd-challenge/esdd-challenges/esdd-2/description, focusing on component-level spoofing, where both speech and environmental sounds may be manipulated or synthesized, creating a more challenging and realistic detection scenario. The challenge will be held in conjunction with the IEEE International Conference on Multimedia and Expo 2026 (ICME 2026) [^4] ICME 2026: https://2026.ieeeicme.org/.
@@ I Introduction
In real-world recording scenarios, audio often contains a mixture of foreground speech and background sounds. We define such audio as comprising two components: (i) speech, i.e., linguistically meaningful speech produced by the primary foreground speaker, and (ii) environmental sound, i.e., any non‑speech background or non‑target speech. With rapid advances in text‑to‑speech, voice conversion, and other generation models, either component can now be modified independently. For example, replacing the background while leaving the foreground human speech unchanged, or modifying the speech content while preserving the background. Such component-level manipulations are harder to detect, as the remaining unaltered component can mislead the systems designed for whole deepfake audio, and they often sound more natural to human listeners.
To address this gap, we introduced a component-level spoofing audio dataset CompSpoof [16] in 2025. Then, we extend the dataset to CompSpoofV2, which contains more than 250,000 audio clips (around 283 hours) formed by mixing bona fide (genuine or real) and spoofed (synthetic or manipulated) audio, including speech and environmental sounds, from multiple sources. We also proposed a separation-enhanced joint learning framework to explore to address this gap.
@@ II Dataset
The CompSpoofV2 is a dataset designed for component-level anti-spoofing detection research, where either the speech or the environmental sound component (or both) may be spoofed. CompSpoofV2 contains over 250,000 audio samples, with a total duration of approximately 283 hours. Each audio sample has a fixed length of 4 seconds and is provided at multiple sampling rates, enabling a more faithful simulation of real-world acoustic and system-level variations. Building upon the CompSpoof dataset[16], CompSpoofV2 significantly expands the diversity of attack sources, environmental sounds, and mixing strategies. In addition, newly generated audio samples are distributed across the test set and are specifically designed to serve as detection data under unseen conditions.
In CompSpoofV2, each audio belongs to one of five categories, covering all combinations of bona fide and spoofed speech and environmental sound as shown in Table I. The task of this challenge is to classify a given audio clip into one of the five defined classes, as shown in Figure 1. The audio sources for each category are summarized in Table II and Table III.
@@ 4 ASV5, MLAAD, New Generated EnvSDD, VcapAV, New Generated
III Evaluation Criteria
@@
In addition, we introduce three auxiliary metrics based on equal error rate (EER): $\mathrm{EER}_{\text{original}}$, $\mathrm{EER}_{\text{speech}}$ and $\mathrm{EER}_{\text{env}}$. $\mathrm{EER}_{\text{original}}$ is the EER for distinguishes the Original ($i=0$) class from all other four classes $i=\{1,2,3,4\}$. Similarly, $\mathrm{EER}_{\text{speech}}$ is the EER for detecting spoofed speech component, while $\mathrm{EER}_{\text{env}}$ is the EER for detecting spoofed environmental component. The three EER metrics are included for diagnostic and analysis purposes only and will not be considered for the leaderboard ranking.
IV Baseline
We propose a separation-enhanced joint learning framework [16] as the baseline for this task. The framework first detects potentially spoofed mixtures, then separates the audio into speech and environmental components, which are then processed by speech- and environment-specific anti-spoofing models. Their outputs are fused and mapped to five-class predictions. The separation and anti-spoofing models are jointly trained to preserve spoofing-relevant cues. The baseline results are as IV.
F1-score Original EER Speech EER Env. EER
Val. set 0.9462 0.0031 0.0172 0.3766
Eval set 0.6224 0.0174 0.1993 0.4336
Test set 0.6327 0.0173 0.1978 0.4279
@@ V Submission Guideline
We use CodaBench platform for results submission [^5] Results Submission: https://www.codabench.org/competitions/12365/. Participants are expected to compute and submit a text file as follows, in which each line contains five columns. From left to right, the columns include:
  • The audio ID;
  • The predicted class ID, which is an integer label from 0 to 4 corresponding to the class definitions in Table I;
  • The confidence score for predicting the original class;
  • The confidence score for the speech component;
  • The confidence score for the environmental sound component.
1 UGBFJ3.wav | 0 | 1.22 | 0.91 | 0.99
2 ABCD12.mp3 | 3 | -1.54 | 0.89 | -1.89
Submitting the scores is not mandatory. If the participant’s method does not generate the scores, they can use a hyphen (-) as a placeholder in the last three columns of the file as example shown in below:
1 UGBFJ3.wav | 0 | - | - | -
2 ABCD12.mp3 | 3 | - | - | -
The challenge has two phases: the preparation phase and the final ranking phase. During the preparation phase, participants submit evaluation set predictions and the leaderboard reflects evaluation set rankings. During the final ranking phase, participants submit test set predictions and the leaderboard determines the final ranking. Each team may submit up to ten results in the final ranking phase, with the best three used for final ranking to ensure fairness. After the competition, we will release the metadata of the evaluation and test sets to support reproducibility and future research.
@@ VI Rules
  • Publicly available pre-trained models released prior to January 1, 2026 are allowed but must be disclosed.
@@
  • The use of any additional synthetic audio–including audio generated by Text-to-Speech models, voice conversion models, or any other audio generation models–is strictly prohibited.
  • Participants may use additional public datasets only with the organizers’ prior approval. Requests must be submitted to the organizers (xueping.zhang@dukekunshan.edu.cn) before February 20, 2026, and the organizers will publish the list of approved datasets; any datasets not on this list are not permitted for training.
@@ VIII Important Dates
Training, validation and evaluation sets release; Baseline code release; Preparation-phase leaderboard update January 30th, 2026
@@
Test set release; Final ranking phase leaderboard update March 20th, 2026